# Assessing High-Risk AI Systems under the EU AI Act: From Legal Requirements to Technical Verification

### v2

@@ 1. Introduction
Progress in AI has created substantial opportunities while also introducing significant technical, organisational, and regulatory challenges, particularly as AI systems are increasingly deployed in critical domains where failures may have serious societal consequences (bengio2024managing; wang2022artificial). In response, regulatory efforts worldwide are intensifying to promote trustworthiness in AI development and deployment (huang2024; smuha2021race; novelli2024taking). Within this context, the EU Artificial Intelligence Act (AI Act) (eu_ai_act_2024) represents one of the most comprehensive regulatory responses to date, establishing a risk-based framework that mandates ex ante verification and ongoing oversight for high-risk AI systems. Despite this progress, a persistent gap remains between high-level legal requirements and the concrete verification activities needed to demonstrate compliance in practice. This gap reflects the absence of explicit and shared mappings between regulatory obligations and implementable assurance practices, resulting in fragmented compliance efforts across stakeholders and Member States and limiting comparability and regulatory learning. In practice, organisations often struggle to translate abstract legal obligations into verifiable controls, tests, and documentation artefacts that can be applied consistently across the AI lifecycle.
These challenges are compounded by three interrelated forms of uncertainty. Interpretive uncertainty concerns the contextual meaning and scope of legal obligations; operational uncertainty relates to how such obligations can be instantiated across heterogeneous development and deployment practices; and procedural uncertainty arises from evolving standards, guidance documents, and the gradual designation of notified bodies (deloitte2024ai; dlapiper2025pause; draghi2024competitiveness; arnal2024implementation; lewis2025regulatorylearning; regulatinguncertainty2025). These uncertainties are further amplified by the emergent behaviour of modern generative AI systems, whose properties arise from complex interactions rather than explicit programming (holtzmanGenerativeModelsComplex2025; koltLessonsComplexSystems2025a). As a horizontal regulation relying on harmonised standards, post-market monitoring, and iterative guidance, the AI Act embodies a model of regulatory learning in which compliance expectations are progressively refined through implementation experience and verification outcomes (zakiConceptualisingOrganisationalPolicy2025; lewis2025regulatorylearning; schrepelAdaptiveRegulation2025). Such learning depends on the availability of structured, comparable, and traceable verification evidence, which remains fragmented across actors, disciplines, and use cases.
@@
Translating these obligations into concrete assessment activities remains challenging. Many provisions of the AI Act are inherently context-dependent, requiring case-specific interpretation grounded in the system’s intended use, operational environment, and evolving risk profile. Requirements related to data governance, robustness, accuracy etc. encompass multiple dimensions, each necessitating distinct validation and testing methodologies, while the lifecycle-oriented nature of the Act entails continuous monitoring and post-deployment evaluation rather than reliance on point-in-time certification alone. Existing initiatives provide partial foundations for AI verification, but none offers a complete and operational mapping between legal requirements and verification activities. The NIST AI Risk Management Framework (NIST_AI_RMF_2025) provides high-level guidance but is not aligned with EU-specific legal obligations, while ISO/IEC standards such as ISO/IEC 42001 (iso_iec_42001_2023) establish process-oriented management requirements without specifying AI Act–specific verification procedures. European standardisation efforts within CEN–CENELEC JTC 21 (CEN-CENELEC_JTC21_2025) aim to develop harmonised standards granting presumption of conformity under Article 40, but this work remains ongoing.
@@
Against this background, the central contribution of this paper is a structured and operational mapping that translates high-level AI Act requirements into concrete verification activities applicable across the AI lifecycle. The mapping is constructed through a systematic decomposition of legal obligations into operational sub-requirements and is grounded in authoritative standards and recognised engineering and governance practices. To ensure consistency and comparability, verification activities are characterised along two dimensions capturing both verification type and lifecycle target. The applicability of the mapping is illustrated through a real-world case study involving a high-risk AI system in the automotive domain. Within this scope, the paper addresses three guiding questions that structure the construction of the proposed mapping: (i) how high-level legal obligations can be decomposed into concrete verification activities; (ii) which verification dimensions are required to ensure comparability and traceability across heterogeneous actors; and (iii) how a shared verification structure can support communication and regulatory learning across use cases and sectors. While related efforts exist (e.g. (hernandez2025open)), the proposed mapping provides a higher level of granularity and operationalisability. It is not intended to be prescriptive or exhaustive, but to serve as a reusable reference that can evolve alongside regulatory guidance, standardisation efforts, and advances in evaluation techniques, thereby supporting more consistent conformity verification and evidence-based regulatory learning over time.
@@ 2.1. Normative inputs and requirement structuring
The starting point of the methodology consists in identifying and structuring a set of high-level requirements derived from the obligations established by the AI Act. Given the heterogeneity and dispersion of these obligations across the regulation, an explicit structuring step is required to organise them into a coherent form suitable for systematic verification. For this purpose, we use the seven principles of Trustworthy AI defined by the European Commission’s High-Level Expert Group in the Ethics Guidelines for Trustworthy AI (eu_trustworthy_ai_2019) as an organising abstraction. Although these principles predate the AI Act and are not legally binding, they have strongly influenced the development of European AI governance and are reflected throughout the regulation, both as explicit obligations (e.g. transparency, human oversight) and as distributed requirements embedded in provisions on governance, documentation, and system performance. Their role in the methodology is therefore structural: they serve as a stable taxonomy for grouping and analysing legally binding requirements, not as an independent source of obligations.
@@ 2.2. Authoritative grounding
To ensure both normative alignment and technical soundness, the operational interpretation of requirements is grounded in authoritative sources that specify or exemplify how high-level obligations are implemented and assessed in practice. These sources are predominantly non-binding, complemented where appropriate by binding instruments whose established operational mechanisms are widely reused in compliance and assurance activities. In this respect, the GDPR (GDPR2016) is included not as a source of normative extension of the AI Act, but because it provides mature and widely operationalised mechanisms—such as risk-based assessment, documentation practices, and accountability procedures—that directly inform how several AI Act obligations are interpreted and verified in practice. The AI Act explicitly operates in complementarity with existing data protection law, and high-risk AI systems in the European context frequently involve personal data processing throughout development, deployment, and monitoring. As a result, verification activities related to data governance, transparency and information provision, record keeping, and affected-person rights often rely on GDPR-derived concepts, artefacts, and procedures, such as information duties, accountability documentation, and impact-assessment practices. For this reason, the GDPR is treated here as an authoritative operational reference within the grounding layer.
@@
Beyond this exception, operational grounding relies on non-binding but authoritative sources that provide reusable mechanisms for implementing and assessing regulatory obligations in practice. These include the EU General Purpose AI Code of Practice (gpai-code-practice-3rd), ISACA’s Advanced Audit in AI Official Review Manual (isaca2025aaia), UNESCO’s Recommendation on the Ethics of AI (unesco2021ethics), the Web Content Accessibility Guidelines (WCAG) 2.2 (wcag22), and a broad range of ISO and IEC standards relevant to AI governance, risk management, quality management, cybersecurity, and data quality (e.g. (ISO_IEC_22989_2022; ISO23894; ISO_IEC_TR_24027_2021; ISO_IEC_24029_2_2023; ISO_IEC_24970_DIS_2025; ISO_IEC_27001_2022; ISO_IEC_27005_2022; ISO_IEC_27035_1_2023; ISO9001; ISO_31000_2018; ISO_IEC_31010_2019; iso_iec_42001_2023; ISO_IEC_5259_1_2024; ISO_8000_8_2015)). These sources were selected because they translate high-level principles into implementable governance processes, controls, and audit practices, thereby bridging normative requirements and operational verification.
@@
Where regulatory, guidance, and standardisation sources did not sufficiently specify technical verification procedures, the analysis was extended to the scientific literature. Highly cited and influential works in AI governance, robustness testing, risk management, and human–system interaction were considered (e.g. (Brundage2020; Basiri2016; Floridi2018; mundhenk2020; sheridan1992telerobotics; schuler1993participatory; CoE_AI_2020)). Selection criteria combined citation impact, venue relevance, and explicit connection to measurable or testable aspects of AI trustworthiness. This layered grounding process ensures that each operational requirement and associated verification activity can be traced back to authoritative legal, standardisation, or scientific sources.
@@ 2.3. Verification dimensions
Concrete verification activities are organised within a structured verification space defined by two orthogonal dimensions: the type of verification and the target of verification within the AI lifecycle. Together, these dimensions provide a common operational language for identifying, organising, and comparing verification activities across heterogeneous systems and stakeholders. The first dimension concerns the type of verification. Two broad categories are distinguished: controls and testing. Controls refer to process-based assurance mechanisms, including governance structures, documentation practices, quality management procedures, and organisational safeguards that ensure responsible system development and operation. Testing refers to empirical evaluations that assess system behaviour, performance, or properties under specified conditions. Controls and testing are complementary rather than hierarchical: controls establish accountability and traceability, while testing provides empirical evidence of compliance. Both may be applied at any stage of the AI lifecycle. The second dimension concerns the target of verification, corresponding to the lifecycle component or artefact being evaluated. Numerous software and AI lifecycle models provide fine-grained decompositions of development and deployment phases (e.g. (de2022artificial; wang2021artificial)). However, such models vary significantly across standards, engineering practices, and organisational contexts, and their level of detail often presupposes technical expertise. For the purpose of compliance verification, which involves regulators, auditors, legal experts, risk managers, and developers alike, a higher level of abstraction is required to support immediate interpretability and shared understanding.
@@ 3.1. R1: Human Agency and Oversight
R1.1: Human in the loop Sheridan et al. (sheridan1992telerobotics) M1.1: Manual override — Confirm that a human operator can take control of the AI system at any time during operation. C P, FP
Sheridan et al. (sheridan1992telerobotics) M1.2: Supervisory control — Assess whether human operators are able to supervise the AI system and intervene when necessary. C P, FP
@@
R1.2: User Autonomy GDPR (Art.7); Andreotta et al. (andreotta2022ai) M1.4: Informed consent — Check that users are provided with clear information about AI usage and are able to give or withhold consent. C FP
@@ 3.2. R2: Technical Robustness and Safety
Basiri et al. (Basiri2016) M2.3: Chaos engineering — Determine whether resilience has been tested through controlled fault injection in production-like environments. T P
@@
R2.2: Robustness to Attacks Brundage et al. (Brundage2020); ISACA manual (§ 2.13.3) M2.5: Red teaming — Examine whether adversarial testing is conducted to uncover vulnerabilities. T P, FP
@@ 3.4. R4: Transparency
Mundhenk et al. (mundhenk2020) M4.2: Saliency maps — Check if visualisation tools highlight relevant input regions. T D, M, FP
@@ 3.5. R5: Diversity, Non-discrimination and Fairness
R5.4: Stakeholder Engagement Schuler et al. (schuler1993participatory) M5.5: Participatory design — Assess whether affected users and communities are involved in design and testing. C FP, P
@@ 3.6. R6: Societal and Environmental Well-being
R6.2: Ethical Alignment Leslie et al. (CoE_AI_2020) M6.2: Rights-based system — Assess whether system design aligns with fundamental rights. C P
@@ 3.7. R7: Accountability
R7.3: Incident Reporting Floridi et al. (Floridi2018) M7.4: Redress systems — Check whether users can report issues or contest AI-driven outcomes. C FP, P
@@ 4. Case Study
This section illustrates how the mapping proposed in Section 3 can be applied in practice through a real, ongoing high-risk AI use case currently being analysed and tested in collaboration with Scania, a Swedish automotive manufacturer. While the mapping is generic and technology-agnostic, this example shows how it can structure assurance activities across the AI lifecycle, from design-time verification to validation and documentation. Table LABEL:tab:use_case presents the results of this mapping process; for the sake of brevity, it is reported only partially and serves as an illustrative example. The use case concerns an AI-based system for detecting cyberattacks within connected vehicles in an industrial setting at Scania. Modern vehicles rely on internal communication networks that enable real-time data exchange between electronic components. Protecting these networks is critical, as cyberattacks may compromise both vehicle safety and system reliability. To address these risks, automotive manufacturers increasingly deploy Intrusion Detection Systems (IDS) that monitor in-vehicle communications and identify abnormal or potentially harmful behaviour. In this setting, the IDS employs AI techniques to analyse in-vehicle network traffic and detect patterns associated with known or emerging attacks. The system processes network data, extracts relevant features, and uses machine-learning models to distinguish normal operation from suspicious activity. To support transparency and engineering oversight, explainability mechanisms are integrated to allow engineers to understand why specific activities are flagged as intrusions. Designed for real-time deployment in vehicle control units or central gateways, this use case demonstrates the application of the proposed mapping in a complex, safety-critical automotive context. The mapping enables structured and traceable verification of data handling, model behaviour, explainability, and documentation, providing evidence to support safety, cybersecurity, and regulatory requirements without intervening in system development.

### v3

@@ 1. Introduction
Progress in AI has created substantial opportunities while also introducing significant technical, organisational, and regulatory challenges, particularly as AI systems are increasingly deployed in critical domains where failures may have serious societal consequences (Bengio et al., 2024; Wang and Chung, 2022). In response, regulatory efforts worldwide are intensifying to promote trustworthiness in AI development and deployment (Huang et al., 2024; Smuha, 2021; Novelli et al., 2024). Within this context, the EU Artificial Intelligence Act (AI Act) (40) represents one of the most comprehensive regulatory responses to date, establishing a risk-based framework that mandates ex ante verification and ongoing oversight for high-risk AI systems. Despite this progress, a persistent gap remains between high-level legal requirements and the concrete verification activities needed to demonstrate compliance in practice. This gap reflects the absence of explicit and shared mappings between regulatory obligations and implementable assurance practices, resulting in fragmented compliance efforts across stakeholders and Member States and limiting comparability and regulatory learning. In practice, organisations often struggle to translate abstract legal obligations into verifiable controls, tests, and documentation artefacts that can be applied consistently across the AI lifecycle.
These challenges are compounded by three interrelated forms of uncertainty. Interpretive uncertainty concerns the contextual meaning and scope of legal obligations; operational uncertainty relates to how such obligations can be instantiated across heterogeneous development and deployment practices; and procedural uncertainty arises from evolving standards, guidance documents, and the gradual designation of notified bodies (Deloitte, 2024; Piper, 2025; Draghi, 2024; Arnal, 2024; Lewis et al., 2025; unknown, 2025). These uncertainties are further amplified by the emergent behaviour of modern generative AI systems, whose properties arise from complex interactions rather than explicit programming (Holtzman et al., 2025; Kolt et al., 2025). As a horizontal regulation relying on harmonised standards, post-market monitoring, and iterative guidance, the AI Act embodies a model of regulatory learning in which compliance expectations are progressively refined through implementation experience and verification outcomes (Zaki, 2025; Lewis et al., 2025; Schrepel, 2025; Mowla and Kabir, 2024). Such learning depends on the availability of structured, comparable, and traceable verification evidence, which remains fragmented across actors, disciplines, and use cases.
@@
Translating these obligations into concrete assessment activities remains challenging. Many provisions of the AI Act are inherently context-dependent, requiring case-specific interpretation grounded in the system’s intended use, operational environment, and evolving risk profile. Requirements related to data governance, robustness, accuracy etc. encompass multiple dimensions, each necessitating distinct validation and testing methodologies, while the lifecycle-oriented nature of the Act entails continuous monitoring and post-deployment evaluation rather than reliance on point-in-time certification alone. Existing initiatives provide partial foundations for AI verification, but none offers a complete and operational mapping between legal requirements and verification activities. The NIST AI Risk Management Framework (National Institute of Standards and Technology (NIST), 2025) provides high-level guidance but is not aligned with EU-specific legal obligations, while ISO/IEC standards such as ISO/IEC 42001 (International Organization for Standardization and International Electrotechnical Commission, 2023) establish process-oriented management requirements without specifying AI Act–specific verification procedures. European standardisation efforts within CEN–CENELEC JTC 21 (CEN-CENELEC Joint Technical Committee 21, 2025) aim to develop harmonised standards granting presumption of conformity under Article 40, but this work remains ongoing.
@@
Against this background, the central contribution of this paper is a structured and operational mapping that translates high-level AI Act requirements into concrete verification activities applicable across the AI lifecycle. The mapping is constructed through a systematic decomposition of legal obligations into operational sub-requirements and is grounded in authoritative standards and recognised engineering and governance practices. To ensure consistency and comparability, verification activities are characterised along two dimensions capturing both verification type and lifecycle target. The applicability of the mapping is illustrated through a real-world case study involving a high-risk AI system in the automotive domain. Within this scope, the paper addresses three guiding questions that structure the construction of the proposed mapping: (i) how high-level legal obligations can be decomposed into concrete verification activities; (ii) which verification dimensions are required to ensure comparability and traceability across heterogeneous actors; and (iii) how a shared verification structure can support communication and regulatory learning across use cases and sectors. While related efforts exist (e.g. (Hernandez et al., 2025)), the proposed mapping provides a higher level of granularity and operationalisability. It is not intended to be prescriptive or exhaustive, but to serve as a reusable reference that can evolve alongside regulatory guidance, standardisation efforts, and advances in evaluation techniques, thereby supporting more consistent conformity verification and evidence-based regulatory learning over time.
@@ 2.1. Normative inputs and requirement structuring
The starting point of the methodology consists in identifying and structuring a set of high-level requirements derived from the obligations established by the AI Act. Given the heterogeneity and dispersion of these obligations across the regulation, an explicit structuring step is required to organise them into a coherent form suitable for systematic verification. For this purpose, we use the seven principles of Trustworthy AI defined by the European Commission’s High-Level Expert Group in the Ethics Guidelines for Trustworthy AI (High-Level Expert Group on Artificial Intelligence, 2019) as an organising abstraction. Although these principles predate the AI Act and are not legally binding, they have strongly influenced the development of European AI governance and are reflected throughout the regulation, both as explicit obligations (e.g. transparency, human oversight) and as distributed requirements embedded in provisions on governance, documentation, and system performance. Their role in the methodology is therefore structural: they serve as a stable taxonomy for grouping and analysing legally binding requirements, not as an independent source of obligations.
@@ 2.2. Authoritative grounding
To ensure both normative alignment and technical soundness, the operational interpretation of requirements is grounded in authoritative sources that specify or exemplify how high-level obligations are implemented and assessed in practice. These sources are predominantly non-binding, complemented where appropriate by binding instruments whose established operational mechanisms are widely reused in compliance and assurance activities. In this respect, the GDPR (European Union, 2016) is included not as a source of normative extension of the AI Act, but because it provides mature and widely operationalised mechanisms—such as risk-based assessment, documentation practices, and accountability procedures—that directly inform how several AI Act obligations are interpreted and verified in practice. The AI Act explicitly operates in complementarity with existing data protection law, and high-risk AI systems in the European context frequently involve personal data processing throughout development, deployment, and monitoring. As a result, verification activities related to data governance, transparency and information provision, record keeping, and affected-person rights often rely on GDPR-derived concepts, artefacts, and procedures, such as information duties, accountability documentation, and impact-assessment practices. For this reason, the GDPR is treated here as an authoritative operational reference within the grounding layer.
@@
Beyond this exception, operational grounding relies on non-binding but authoritative sources that provide reusable mechanisms for implementing and assessing regulatory obligations in practice. These include the EU General Purpose AI Code of Practice (Chairs and Vice‑Chairs of the General‑Purpose AI Code of Practice, 2025), ISACA’s Advanced Audit in AI Official Review Manual (ISACA, 2025), UNESCO’s Recommendation on the Ethics of AI (UNESCO, 2021), the Web Content Accessibility Guidelines (WCAG) 2.2 (World Wide Web Consortium (W3C), 2023), and a broad range of ISO and IEC standards relevant to AI governance, risk management, quality management, cybersecurity, and data quality (e.g. (27; 26; 28; 5; 3; 24; 25; 23; 31; 41; 42; International Organization for Standardization and International Electrotechnical Commission (2023); 4; 11)). These sources were selected because they translate high-level principles into implementable governance processes, controls, and audit practices, thereby bridging normative requirements and operational verification.
@@
Where regulatory, guidance, and standardisation sources did not sufficiently specify technical verification procedures, the analysis was extended to the scientific literature. Highly cited and influential works in AI governance, robustness testing, risk management, and human–system interaction were considered (e.g. (Brundage and et al., 2020; Basiri et al., 2016; Floridi et al., 2018; Mundhenk et al., 2020; Sheridan, 1992; Schuler and Namioka, 1993; Leslie et al., 2020)). Selection criteria combined citation impact, venue relevance, and explicit connection to measurable or testable aspects of AI trustworthiness. This layered grounding process ensures that each operational requirement and associated verification activity can be traced back to authoritative legal, standardisation, or scientific sources.
@@ 2.3. Verification dimensions
Concrete verification activities are organised within a structured verification space defined by two orthogonal dimensions: the type of verification and the target of verification within the AI lifecycle. Together, these dimensions provide a common operational language for identifying, organising, and comparing verification activities across heterogeneous systems and stakeholders. The first dimension concerns the type of verification. Two broad categories are distinguished: controls and testing. Controls refer to process-based assurance mechanisms, including governance structures, documentation practices, quality management procedures, and organisational safeguards that ensure responsible system development and operation. Testing refers to empirical evaluations that assess system behaviour, performance, or properties under specified conditions. Controls and testing are complementary rather than hierarchical: controls establish accountability and traceability, while testing provides empirical evidence of compliance. Both may be applied at any stage of the AI lifecycle. The second dimension concerns the target of verification, corresponding to the lifecycle component or artefact being evaluated. Numerous software and AI lifecycle models provide fine-grained decompositions of development and deployment phases (e.g. (De Silva and Alahakoon, 2022; Wang et al., 2021)). However, such models vary significantly across standards, engineering practices, and organisational contexts, and their level of detail often presupposes technical expertise. For the purpose of compliance verification, which involves regulators, auditors, legal experts, risk managers, and developers alike, a higher level of abstraction is required to support immediate interpretability and shared understanding.
@@ 3.1. R1: Human Agency and Oversight
R1.1: Human in the loop Sheridan et al. (Sheridan, 1992) M1.1: Manual override — Confirm that a human operator can take control of the AI system at any time during operation. C P, FP
Sheridan et al. (Sheridan, 1992) M1.2: Supervisory control — Assess whether human operators are able to supervise the AI system and intervene when necessary. C P, FP
@@
R1.2: User Autonomy GDPR (Art.7); Andreotta et al. (Andreotta et al., 2022) M1.4: Informed consent — Check that users are provided with clear information about AI usage and are able to give or withhold consent. C FP
@@ 3.2. R2: Technical Robustness and Safety
Basiri et al. (Basiri et al., 2016) M2.3: Chaos engineering — Determine whether resilience has been tested through controlled fault injection in production-like environments. T P
@@
R2.2: Robustness to Attacks Brundage et al. (Brundage and et al., 2020); ISACA manual (§ 2.13.3) M2.5: Red teaming — Examine whether adversarial testing is conducted to uncover vulnerabilities. T P, FP
@@ 3.4. R4: Transparency
Mundhenk et al. (Mundhenk et al., 2020) M4.2: Saliency maps — Check if visualisation tools highlight relevant input regions. T D, M, FP
@@ 3.5. R5: Diversity, Non-discrimination and Fairness
R5.4: Stakeholder Engagement Schuler et al. (Schuler and Namioka, 1993) M5.5: Participatory design — Assess whether affected users and communities are involved in design and testing. C FP, P
@@ 3.6. R6: Societal and Environmental Well-being
R6.2: Ethical Alignment Leslie et al. (Leslie et al., 2020) M6.2: Rights-based system — Assess whether system design aligns with fundamental rights. C P
@@ 3.7. R7: Accountability
R7.3: Incident Reporting Floridi et al. (Floridi et al., 2018) M7.4: Redress systems — Check whether users can report issues or contest AI-driven outcomes. C FP, P
@@ 4. Case Study
This section illustrates how the mapping proposed in Section 3 can be applied in practice through a real, ongoing high-risk AI use case currently being analysed and tested in collaboration with Scania, a Swedish automotive manufacturer. While the mapping is generic and technology-agnostic, this example shows how it can structure assurance activities across the AI lifecycle, from design-time verification to validation and documentation. Table LABEL:tab:use_case presents the results of this mapping process. The use case concerns an AI-based system for detecting cyberattacks within connected vehicles in an industrial setting at Scania. Modern vehicles rely on internal communication networks that enable real-time data exchange between electronic components. Protecting these networks is critical, as cyberattacks may compromise both vehicle safety and system reliability. To address these risks, automotive manufacturers increasingly deploy Intrusion Detection Systems (IDS) that monitor in-vehicle communications and identify abnormal or potentially harmful behaviour. In this setting, the IDS employs AI techniques to analyse in-vehicle network traffic and detect patterns associated with known or emerging attacks. The system processes network data, extracts relevant features, and uses machine-learning models to distinguish normal operation from suspicious activity. To support transparency and engineering oversight, explainability mechanisms are integrated to allow engineers to understand why specific activities are flagged as intrusions. Designed for real-time deployment in vehicle control units or central gateways, this use case demonstrates the application of the proposed mapping in a complex, safety-critical automotive context. The mapping enables structured and traceable verification of data handling, model behaviour, explainability, and documentation, providing evidence to support safety, cybersecurity, and regulatory requirements without intervening in system development.
Table 12. Results of the mapping instantiation for the Scania use case.